# FineVQ: Fine-Grained User Generated Content Video Quality Assessment

### v1

@@ Abstract
The rapid growth of user-generated content (UGC) videos has produced an urgent need for effective video quality assessment (VQA) algorithms to monitor video quality and guide optimization and recommendation procedures. However, current VQA models generally only give an overall rating for a UGC video, which lacks fine-grained labels for serving video processing and recommendation applications. To address the challenges and promote the development of UGC videos, we establish the first large-scale Fine-grained Video quality assessment Database, termed FineVD, which comprises 6104 UGC videos with fine-grained quality scores and descriptions across multiple dimensions. Based on this database, we propose a Fine-grained Video Quality assessment (FineVQ) model to learn the fine-grained quality of UGC videos, with the capabilities of quality rating, quality scoring, and quality attribution. Extensive experimental results demonstrate that our proposed FineVQ can produce fine-grained video-quality results and achieve state-of-the-art performance on FineVD and other commonly used UGC-VQA datasets. Both FineVD and FineVQ will be made publicly available.
@@ 1 Introduction
Recent years have witnessed an explosion of user-generated content (UGC) videos [36, 75], thanks to the evolution of video processing technologies, and the popularity of social media platforms [52, 71]. UGC videos cover diverse video capturing and processing conditions [34], thus commonly suffering from various degradations such as noise, blur, shaking, etc., which makes the visual quality of UGC videos vary greatly. Understanding and predicting the quality of UGC videos has long been an important but unsolved problem, which can be applied to capture devices and social media sites to monitor video quality and guide optimization and recommendation procedures [71]. However, different applications generally require quality labels considered from different dimensions, making a single overall quality rating insufficient for diverse downstream tasks [12].
@@
Based on the FineVD benchmark, we further introduce the first Fine-grained Video Quality evaluator (FineVQ) to perform multi-dimensional fine-grained video quality assessment in a one-for-all manner. To achieve the comprehensive capabilities of quality rating, quality scoring, and quality depicting from multiple dimensions, FineVQ is built upon large multimodal models (LMMs) [25, 6, 70] and leverages instruction tuning [31] and low-rank adaptation (LoRA) [18] techniques. Specifically, our FineVQ adopts the vision encoder used in InternVL [6], i.e., InternViT, as the spatial visual backbone, along with a motion feature extraction encoder [13] as the temporal visual backbone. A large language model, i.e., IternLM [50], is used as the language backbone to facilitate the quality-related score regression and text generation. Moreover, to extract quality-related features and refine quality-specific attributes, we impose LoRA weights on both the spatial visual backbone and the large language backbone. Additionally, we also employ instruction tuning by inputting various question-answering pairs and training the feature projectors to enable task-specific quality awareness. Extensive experimental results demonstrate the fine-grained quality rating, scoring, depicting abilities, and the state-of-the-art performance of the FineVQ. The main highlights of this work include:
@@ 3.2 In-lab Subjective Study
Based on the collected quality scores and attribute labels, we further prepare language question-answering (QA) pairs for tuning LMMs [65, 73, 74]. Specifically, we first generate QA pairs automatically using an excellent large language model, i.e., GPT-4 [1], and then we revise the generated texts manually to avoid ambiguous QA pairs. Specifically, we introduce three tasks including quality-level prediction, quality score regression, and quality description, respectively, as shown in Figure 2(c). The generated QA pairs can significantly facilitate the subsequent training for our FineVQ model.
@@ 3.3 Subjective Data Analysis
We follow the suggestions given in [45] to conduct the outlier detection and subject rejection [11]. As a result, no subject is rejected, and each image is rated by 22 valid subjects. Among all scores, about 3.16% of the total subjective evaluations are identified as outliers and are subsequently removed. For the remaining valid subjective ratings, we convert these raw ratings into Z-scores as follows:
$z_{ij}=\frac{m_{ij}-\mu_{i}}{\sigma_{i}},$ (1)
@@
$\text{MOS}_{j}=\frac{1}{N}\sum_{i=1}^{N}z_{ij}^{\prime},\vspace{-6pt}$ (2)
@@ 4.3 Instruction Tuning and LoRA Adaptation
$\textbf{h}=\textbf{W}\textbf{x}+\Delta\textbf{W}\textbf{x}=(\textbf{W}+\textbf{A}\textbf{B})\textbf{x}.$ (5)
@@ 5.2 Evaluation on the FineVD
For the quality attribute prediction task, we use the same dataset split setting as mentioned above. A total of 7 large multimodal models that accept video as input are selected as benchmark methods, including VideoLLaMA2 [7], Video-LLaVA [29], VideoChat2 [26], Video-ChatGPT [35], InternVL2 [6], MiniCPM-V [69], Qwen2-VL [55], LLaVA-NeXT [22]. We set two types of questions including “yes-or-no” and “which” to test the ability of LMMs and our FineVQ on quality-related low-level vision tasks [63]. Specifically, for “yes-or-no” questions, we infer the LMMs with the prompts of “Is these any $<$dimension$>$ distortion in this video?”, while for “which” questions, we infer the LMMs with the prompts of “Which distortion exists in this video?” and “Which distortion has the most impact on the quality of this video?” We adopt prediction accuracy as the metric for evaluating the performance of the quality attribute prediction task.
@@
Results on The Quality Attribute Prediction Task.

### v2

@@ Abstract
The rapid growth of user-generated content (UGC) videos has produced an urgent need for effective video quality assessment (VQA) algorithms to monitor video quality and guide optimization and recommendation procedures. However, current VQA models generally only give an overall rating for a UGC video, which lacks fine-grained labels for serving video processing and recommendation applications. To address the challenges and promote the development of UGC videos, we establish the first large-scale Fine-grained Video quality assessment Database, termed FineVD, which comprises 6104 UGC videos with fine-grained quality scores and descriptions across multiple dimensions. Based on this database, we propose a Fine-grained Video Quality assessment (FineVQ) model to learn the fine-grained quality of UGC videos, with the capabilities of quality rating, quality scoring, and quality attribution. Extensive experimental results demonstrate that our proposed FineVQ can produce fine-grained video-quality results and achieve state-of-the-art performance on FineVD and other commonly used UGC-VQA datasets. Both FineVD and FineVQ are publicly available at: https://github.com/IntMeGroup/FineVQ.
@@ 1 Introduction
Recent years have witnessed an explosion of user-generated content (UGC) videos [37, 76], thanks to the evolution of video processing technologies, and the popularity of social media platforms [53, 72]. UGC videos cover diverse video capturing and processing conditions [35], thus commonly suffering from various degradations such as noise, blur, shaking, etc., which makes the visual quality of UGC videos vary greatly. Understanding and predicting the quality of UGC videos has long been an important but unsolved problem, which can be applied to capture devices and social media sites to monitor video quality and guide optimization and recommendation procedures [72]. However, different applications generally require quality labels considered from different dimensions, making a single overall quality rating insufficient for diverse downstream tasks [13, 12].
@@
Based on the FineVD benchmark, we further introduce the first Fine-grained Video Quality evaluator (FineVQ) to perform multi-dimensional fine-grained video quality assessment in a one-for-all manner. To achieve the comprehensive capabilities of quality rating, quality scoring, and quality depicting from multiple dimensions, FineVQ is built upon large multimodal models (LMMs) [26, 6, 71] and leverages instruction tuning [32] and low-rank adaptation (LoRA) [19] techniques. Specifically, our FineVQ adopts the vision encoder used in InternVL [6], i.e., InternViT, as the spatial visual backbone, along with a motion feature extraction encoder [14] as the temporal visual backbone. A large language model, i.e., InternLM [51], is used as the language backbone to facilitate the quality-related score regression and text generation. Moreover, to extract quality-related features and refine quality-specific attributes, we impose LoRA weights on both the spatial visual backbone and the large language backbone. Additionally, we also employ instruction tuning by inputting various question-answering pairs and training the feature projectors to enable task-specific quality awareness. Extensive experimental results demonstrate the fine-grained quality rating, scoring, depicting abilities, and the state-of-the-art performance of FineVQ. The main highlights of this work include:
@@ 3.2 In-lab Subjective Study
Based on the collected quality scores and attribute labels, we further prepare language question-answering (QA) pairs for tuning LMMs [66, 74, 75]. Specifically, we first generate QA pairs automatically using an excellent large language model, i.e., GPT-4 [1], and then we revise the generated texts manually to avoid ambiguous QA pairs. Finally, we introduce three tasks including quality-level prediction, quality score regression, and quality description, respectively, as shown in Figure 2(c). The generated QA pairs can significantly facilitate the subsequent training for our FineVQ model.
@@ 3.3 Subjective Data Analysis
We follow the suggestions given in [46] to conduct the outlier detection and subject rejection. As a result, no subject is rejected, and each image is rated by 22 valid subjects. Among all scores, about 3.16% of the total subjective evaluations are identified as outliers and are subsequently removed. For the remaining valid subjective ratings, we convert these raw ratings into Z-scores [11] as follows:
$z_{ij}=\frac{m_{ij}-\mu_{i}}{\sigma_{i}},\vspace{-0.5em}$ (1)
@@
$\text{MOS}_{j}=\frac{1}{N}\sum_{i=1}^{N}z_{ij}^{\prime},\vspace{-0.5em}$ (2)
@@ 4.3 Instruction Tuning and LoRA Adaptation
$\textbf{h}=\textbf{W}\textbf{x}+\Delta\textbf{W}\textbf{x}=(\textbf{W}+\textbf{A}\textbf{B})\textbf{x}.\vspace{-0.5em}$ (5)
@@ 5.2 Evaluation on the FineVD
For the quality attribute prediction task, we use the same dataset split setting as mentioned above. A total of 7 large multimodal models that accept video as input are selected as benchmark methods, including VideoLLaMA2 [7], Video-LLaVA [30], VideoChat2 [27], Video-ChatGPT [36], InternVL2 [6], MiniCPM-V [70], Qwen2-VL [56], LLaVA-NeXT [23]. We set two types of questions including “yes-or-no” and “which” to test the ability of LMMs and our FineVQ on quality-related low-level vision tasks [64]. Specifically, for “yes-or-no” questions, we query the LMMs with the prompts of “Is these any $<$dimension$>$ distortion in this video?”, while for “which” questions, we infer the LMMs with the prompts of “Which distortion exists in this video?” and “Which distortion has the most impact on the quality of this video?” We adopt prediction accuracy as the metric for evaluating the performance of the quality attribute prediction task.
@@
Table 7: Running time comparison with other methods.
Model NIQE [39] QAC [69] HOSA [68] TLVQM [22] VIDEVAL [53] RAPIQUE [54] VSFA [24] GSTVQA [2] SimpleVQA [49] FastVQA [60] Dover [62] FineVQ
Time(s)$\downarrow$ 918.4 369.0 250.0 848.2 1881 152.7 95.05 164.1 71.62 13.12 9.006 18.57
[…]
@@ 5.5 Ablation Study
Runtime.
We also compare the runtime and FPS of state-of-the-art models on 10 2K 240-frame videos with an A6000 card or matlab and show results in Table 7. FineVQ achieves comparable results.
@@ 6 Conclusion
Acknowledgment
This work was supported in part by the NSFC 62401365, 62225112, 62271312, 62132006, 62271308, in part by the STCSM 24ZR1432000, 24511106902, 22511105700, 22DZ2229005, and in part by the 111 plan BP0719010.